Suppose the production function is given by Q (L,k)=L 3\2 k1\2 assuming capital fixed find APL and MPL
1
Expert's answer
2020-06-08T11:06:38-0400
"AP_L=\frac{TP}{L}"


"AP_L=\frac{L^{\frac{3}{2}} K^{\frac{1}{2}}}{L}=(LK)^{\frac{1}{2}}"


"MP_L=\frac{\varDelta TP_L}{ \varDelta L}"


"MP_L=\frac{3}{2}L^{\frac{1}{2}}K^{\frac{1}{2}}"